Zeronknight (zeronknight):

Build and solve the following inequality: * Ten less than the product of a number and negative four is less than or equal to negative six. x ≤ - 4 x ≥ - 1 x ≤ 16 x ≥ 4

Zeronknight (zeronknight):

@Luigi0210

OpenStudy (luigi0210):

Have you tried writing it out?

Zeronknight (zeronknight):

I don't know how to even write this out, I was never good at word problems with inequalities.

Zeronknight (zeronknight):

All I know is there is <_ a 10 and a -4

OpenStudy (luigi0210):

At least you understand some of the wording (: "less than" meaning subtraction "product of" meaning multiplication "less than or equal to" meaning \(\le\) Putting them together: \(\large -4x-10 \le -6 \)

OpenStudy (luigi0210):

Think you can solve it from here?

Zeronknight (zeronknight):

add 10 to both sides so 10 cancels out 4 turns into 6 and 6 turns in 4

Zeronknight (zeronknight):

Do i need to divide?

OpenStudy (luigi0210):

Oh, and when dealing with "less than," you have to reverse the order of the numbers and correct, now you're just left with \(\large -4x \le 4 \) and to get x alone, yes, you'll need to divide.

Zeronknight (zeronknight):

so 4 / 4 = 1 But since 4 is negative 1 becomes negative .

Zeronknight (zeronknight):

We need to flip the sign again, right?

OpenStudy (luigi0210):

right (:

Zeronknight (zeronknight):

So X >_ 1

Zeronknight (zeronknight):

-1*

OpenStudy (luigi0210):

And you got bingo :D

Zeronknight (zeronknight):

Alright, thanks for your help! C: